Two cans of cola and one can of root beer contain 149 grams of sugar. Three
cans of cola and four cans of root beer contain 336 grams of sugar. How many
grams of sugar are there in each?

Answer by Cromlix(4381)

Hi there,
Cola = 'x'
root beer = 'y'
2x + y = 149......(1)
3x + 4y = 336.....(2)
Multiply (1) by 4
8x + 4y = 596.....(1)
3x + 4y = 336.....(2)
Subtract (2) from (1)
5x = 260
x = 52
Substitute x = 52 in (1)
2(52) + y = 149
104 + y = 149
y = 149 - 104
y = 45
Cola has 52g of sugar in it.
Root Beer has 45g of sugar in it.
Hope this helps :-)